OpenStudy (anonymous):

Patricia has a lawn, ABCD. She has placed a watering hose, BD, as shown below.

OpenStudy (anonymous):

Part A: Patricia plans to put a fence along the length AD of her lawn. What is the length of the fence required? Part B: Using complete sentences, explain how you arrived at the answer for Part A.

OpenStudy (anonymous):

please help! i dont remember how to solve this and its really frustrating

OpenStudy (raden):

first, u have to find out the length of a watering hose, BD use the pythagorean theorem : BD^2 = BC^2 + CD^2

OpenStudy (anonymous):

Okay so this would be BD^2= 30^2 +16^2? which should equal to 900+256= 1156.. but you still need to square it.. which is 1336336..so this means that would be equal to BD? that doesnt sound right lol i must have done somthing wrong

OpenStudy (raden):

BD^2 = BC^2 + CD^2 BD^2 = (30)^2 + (16)^2 BD^2 = 900 + 256 BD^2 = 1156 take the squareroot, BD = sqrt(1156) BD = 34

OpenStudy (anonymous):

okay was close, so now, what must i do with the measure of BD? it says "Patricia plans to put a fence along the length AD of her lawn. What is the length of the fence required? am i supposed to find the meansure of AD too?

OpenStudy (raden):

yes, then look at the triangle of ABD, are u familiar with the 30-60-90 right triangle ?

OpenStudy (anonymous):

yes i am

OpenStudy (raden):